There’s a big difference between

∫ ∫

and AMS-TEX’s

∫∫

AMS-TEX also offers

∫∫∫ , ∫∫∫∫ , ∮ , and ∫ ⋯ ∫

AMS-TEX offers superior typesetting of mathematics.

∫∫∫_{M} dω = ∫∫_{∂M} ω

∫_{∂(M−∪_{i=1}^{n}U_{i})} r^{∗}Π = ∑_{i=1}^{n} ∫_{∂U_{i}} r^{∗}Π

∑_{\begin{matrix}0≤i≤m \\ 0≤j≤n \end{matrix}} P (i, j)